OpenStudy (anonymous):

Suppose f(π/3) = 2 and f '(π/3) = −3, and let g(x) = f(x) sin x and h(x) = (cos x)/f(x). Find the following

hero (hero):

Find what?

OpenStudy (anonymous):

i bet it is \[g'(\frac{\pi}{3})\]

OpenStudy (anonymous):

yes lol

hero (hero):

More proof that satellite needs to retire

OpenStudy (anonymous):

by the product rule \[g'(x)=f'(x)\sin(x)+f(x)\cos(x)\] now plug in \[\frac{\pi}{3}\] so get your answer

hero (hero):

When you know the question before it is even asked, it's time to retire.

OpenStudy (anonymous):

what else could it be??

OpenStudy (anonymous):

i did all of that and i keep getting a wrong answer

OpenStudy (anonymous):

same for next one but use the quotient rule. really? lets try it

OpenStudy (anonymous):

\[g'(x)=f'(x)\sin(x)+f(x)\cos(x)\] \[g'(\frac{\pi}{3})=f'(\frac{\pi}{3})\sin(\frac{\pi}{3})+f(\frac{\pi}{3})\cos(\frac{\pi}{3})\] \[g'(x)=-3\times \frac{\sqrt{3}}{2}+2\times \frac{1}{2}\] is as start

OpenStudy (anonymous):

we get \[1-\frac{3\sqrt{3}}{2}\] or if you prefer \[\frac{2-3\sqrt{3}}{2}\]

OpenStudy (anonymous):

i did that too : ))) but its telling me its a wrong answer lol

OpenStudy (anonymous):

and got the same answer as you...

OpenStudy (anonymous):

then be happy with your answer and tell "it" to @#$% off

OpenStudy (anonymous):

maybe a syntax error? i take it this is on line homework

OpenStudy (anonymous):

yes its a webassign crap and i am so mad.... thank you for your help

OpenStudy (anonymous):

i am familiar with web assign. best bet is to contact your teacher and say "this is what i think is right"

OpenStudy (anonymous):

okay thank a lot : )))

OpenStudy (anonymous):

yw